Hello i really need help on a wriiten equation im not good on those it says open ended , suppose you used tue distrubative property to get the expression 3m-6n-15 ,with what expression could you have started ?
Answer by nyc_function(2741)

If you were given 3m-6n-15 and want to find the original expression, just factor out a common term and then divide it by every term found in the expression 3m-6n-15.
The common factor is 3.
3m/3 = m
-6n/3 = -2n
-15/3 = -5
You started with 3(m - 2n - 5).